Calculate de broglie wavelength and velocity of an electron carrying kinetic energy 3keV

1
Expert's answer
2022-02-20T15:44:54-0500

We find the speed of an electron, knowing its kinetic energy"W_k=\frac{m_e \cdot v^2}{2}"

"v=\sqrt{\frac{2 \cdot W_k}{m_e}}=\sqrt{\frac{2 \cdot 3 \cdot 10^3 \cdot 1.6 \cdot 10^{-19}}{9.1 \cdot 10^{-31}}}=3.248 \cdot 10^7 m/s"

the speed of an electron is much less than the speed of light

"v<< c"

therefore, finds the de Broglie wavelength by the formula

"\lambda_{db}=\frac{h}{m_e \cdot v}= \frac{6.62 \cdot 10^{-34}}{9.1 \cdot 10^{-31}\cdot 3.248 \cdot 10^7 }=2.24 \cdot 10^{-11}m"